# Vertex-Minimal Paper Tori

### v4

@@ 1.1 Context and Main Results
A flat torus is a quotient of the form $\mbox{\boldmath{$R$}}^{2}/\Lambda$, where $\Lambda$ is a lattice of translations of $\mbox{\boldmath{$R$}}^{2}$. More concretely, a flat torus is what you get when you identify the opposite sides of a parallelogram by translations.
A polyhedral torus is a continuous, piecewise affine embedding $\phi:T\to\Omega\subset\mbox{\boldmath{$R$}}^{3}.$ Here $T$ is a flat torus that has been triangulated, and $\phi$ is a continuous embedding that is affine on each triangle of the triangulation. If, additionally, $\phi$ is an affine isometry on each triangle, we say that $\phi$ is flat. When $\phi$ is a flat embedding, $\phi$ tells how to build a torus in $\mbox{\boldmath{$R$}}^{3}$ out of finitely many triangles so that the cone angle around each vertex is $2\pi$. We call a flat embedded polyhedral torus a paper torus. Some authors would call this an origami torus.
Surprisingly, paper tori exist. The 1960 paper of Y. Burago and V. Zalgaller [BZ1] gives the first constructions. The 1985 paper [BZ2] proves that one can realize every isometry class of flat torus as a paper torus. Their construction produces examples with thousands of faces. The 2024 preprint of F. Lazarus and F. Tallerie [LT] gives a universal combinatorial type of triangulation which does the job simultaneously for all isometry types.
The works of P. Arnoux, S. Lelievre, and A. Malaga [ALM] and (independently) T. Tsuboi [T] give a different construction which achieves every isometry class of flat torus as a paper torus, using far fewer faces than in the other cited works. With minor differences in the final results, both papers essentially prove that the paper tori described by U. Brehm [Br] in 1978 achieve every isometry class of flat torus. In 2025, Vincent Tugayé [Tu], discovered a $9$-vertex paper torus.
One needs at least $7$ vertices to make a paper torus because there is no $6$ vertex triangulation of a torus. There is a $7$-vertex triangulation of a torus, and it is unique up to combinatorial isomorphism. This triangulation, sometimes called the Moebius triangulation, is shown in Figure 1.1 below. The Moebius triangulation gives the famous embedding of the complete graph $K_{7}$ in a torus. Figure 1.1 shows part of the universal cover of the Moebius torus, as well as a fundamental domain for it.
@@
The example in Theorem 1.2 has all $8$ points in its convex hull boundary and $2$-fold symmetry. We will discuss it below in great detail.
1.2 Seven Vertices: Proof Sketch
@@
In order to keep this paper self-contained, I will give a light and entirely combinatorial computer-assisted proof of the Hull Lemma. I will also derive the Hull Theorem from the Hull Lemma, though this is not needed for our results.
1.3 Eight Vertices: Experimentation
@@
There are $7$ combinatorial types of triangulation of the torus having $8$ vertices. Of these $7$ there is (in my opinion) a worst triangulation and a best triangulation. The worst triangulation is obtained from the Moebius triangulation by subdividing one of the triangles. The best triangulation has degree sequence $66666666$ and is vertex-transitive. This triangulation is very much like the Moebius triangulation. Figure 1.2 below shows part of the universal cover of this triangulation, as well as a fundamental domain. I will explain the meaning of the blue triangles below.
For all but the worst triangulation, I generated lots of embedded examples by randomly sampling collections of $8$ points on the unit sphere and testing the embedding condition.
@@
Next, I ran a hill-climbing algorithm which started with random embedded examples on the unit sphere and repeatedly perturbed them – not necessarily keeping them on the unit sphere – with the goal of minimizing an objective flatness function. The flatness function computes the maximum deviation of a cone angle from being $2\pi$. After a lot of experimentation, I implemented $4$ tricks that helped immensely.
@@
  3. 3.
  4. 4.
    It helped to force the examples to have $2$-fold rotational symmetry. This simplification came later. After I found some asymmetric examples, Samuel Lelievre asked me about symmetric examples and I tweaked the program. The program works more reliably after incorporating the symmetry.
1.4 Meet the Pup Tent
With all the tricks above in place, and a lot of playing around, I found some embedded and $2$-fold symmetric examples of face number $6$ which are flat up to $10^{-16}$. I took the nicest one and simplified the coordinates, feeding it back into the program to make it near-flat again. Then I implemented a high precision version of Newton’s method and improved my example to one that was flat up to $10^{-400}$. Using a different method, Peter Doyle computed the first million digits of my example!
Here are the coordinates for my example, truncated after $32$ digits. Theorem 1.2 asserts more specifically that this example is within $10^{-32}$ of an $8$-vertex paper torus.
@@ 1.4 Meet the Pup Tent
The 0th row of this array gives coordinates for vertex 0, and so on. The example has $2$-fold symmetry which reflects in the $Z$-axis and implements the vertex permutation $(04)(13)(26)(57)$. You can operate my software and generate up to about $1000$ digits of precision.
Based on his computer plots, Doyle suggested that I call this example the pup tent. Let us say more formally that a pup tent is an $8$-vertex paper torus with $2$-fold rotational symmetry, based on the best triangulation, and having $6$ faces in the convex hull boundary arranged combinatorially like the blue triangles in Figure 1.2.
Figure 1.3: The intrinsic structure.
Figure 1.3 shows a numerical computation of the intrinsic geometric structure of the faces of the pup tent. We are showing part of the universal cover of the triangulation. You can generate your own version of Figure 1.3 using the following $8$ vertices and the two generators $\lambda_{1},\lambda_{2}$ of the deck group.
[…]
Figure 1.3 looks quite tame, but the extrinsic geometry of the pup tent is beautiful and very intricate. Figure 1.4 shows one of the slices of the pup tent by a plane parallel to the $Z$ axis. The solid handlebody bounded by the torus is colored orange. This slice is very spiky and sort of “barely embedded”, but the embeddedness is robust compared to the plot precision.
Figure 1.4: A horizontal slice of the pup tent
Figure 1.5 shows the slice by the $XZ$ plane, which goes through the axis of symmetry of the torus. The slice of the handlebody is an annulus. The two “parallel vertical segments” on top are actually each unions of two segments which meet in the middle at angle that is close to $\pi-.01$. We show a closeup of the top central area to highlight the fact that the two polygons really are disjoint.
Figure 1.5: The $XZ$ planar slice of the pup tent.
Figure 1.6 shows the slice by the $YZ$ plane, with some closeups.
Figure 1.6: The $YZ$ planar slice of the pup tent
The left side of Figure 1.7 shows a random slice of the pup tent.
Figure 1.7: Extrinsic and Intrinsic structures compared
In §6, I include Doyle’s plots of the pup tent. I also include a sheet of paper I made, after consulting with Alba Malaga and Noah Mongtomery, which you can fold up into your own example.
1.5 Eight Vertices: Proof Sketch
If we leave the variables $z_{0},z_{1},z_{2}$ unspecified, the example in Equation 1 is part of a $3$-parameter family of examples we call $T(z_{0},z_{1},z_{2})$. Let $\theta_{k}(z_{0},z_{1},z_{2})$ be the cone angle of $T(z_{0},z_{1},z_{2})$ at vertex $k$. Let
$F(z_{0},z_{1},z_{2})=(\theta_{0},\theta_{1},\theta_{2}).$
If $F(z_{0},z_{1},z_{2})=(2\pi,2\pi,2\pi)$ and $T(z_{0},z_{1},z_{2})$ is embedded then we have our proof. We use the Inverse Function Theorem. Let $dF$ be the Jacobian of $F$, computed at our example. We display $dF$ and its inverse up to $2$ digits:
$\left[\begin{matrix}-0.91\ldots&+0.74\ldots&+0.39\ldots\\ +0.74\ldots&-1.92\ldots&+1.14\ldots\\ +0.39\ldots&+1.14\ldots&-0.06\ldots\end{matrix}\right]\qquad\left[\begin{matrix}-0.56\ldots&+0.23\ldots&+0.76\ldots\\ +0.23\ldots&-0.04\ldots&+0.64\ldots\\ +0.76\ldots&+0.64\ldots&+0.57\ldots\end{matrix}\right]$ (3)
Notice that $dF$ is very well conditioned because $dF^{-1}$ is small. Our proof is an interplay of three things. First, we show that our example is robustly embedded in the sense that moderate perturbations keep it embedded. Second, we show that our example is very flat. Third we show, by bounding derivatives, that our Jacobian matrix stays fairly constant and well conditioned through a healthy neighborhood of our example. We then use an effective version of the Inverse Function Theorem so show that a true paper torus is quite near our example. In many parts of the proof, we work with a scale-up version of our example so that we can work with integer calculations.
Peter Doyle tells me that he very probably can show that my example is near a flat example by using the method in Matthew Ellison’s 2023 preprint [E].
Here is a quick corollary of our method.
Corollary 1.4
There exists a $6$-dimensional open manifold of pup tents.
Proof: We can normalize the vertices we work with so that the line of symmetry is the $Z$-axis. We can then perturb our points symmetrically by changing the $X$ and $Y$ coordinates but not the $Z$-coordinates. As long as the perturbation is small enough, the Inverse Function Theorem selects $Z$-coordinates for us which give a pup tent. We have $8$ free parameters to play with and then these are grouped into a $2$-dimensional foliation under the action of similarities of the $XY$-plane which fix the origin. A cross-section of this foliation gives us our $6$-dimensional open manifold of inequivalent pup tents. $\spadesuit$
1.6 Paper Organization
  • In §2, we will prove Theorem 1.1. At the end of §2 I derive the Hull Theorem from the Hull Lemma. This is not needed for any of the results in the paper.
  • In §3, we prove that our example $T$ from Equation 1 is $10^{-4}$-robustly embedded, meaning that if we change any of $z_{0},z_{1},z_{2}$ by less than $10^{-4}$ then the result is still embedded.
  • In §4, we prove that $T$ is $10^{-30}$-flat, meaning that the deviation of each cone angle from $2\pi$ is less than $10^{-30}$
  • In §5, we use the results from §3-4 and finish the proof of Theorem 1.2.
  • in §6, I include some additional pictures and some other information.
1.7 Computer Assistance
My proof is heavily computer assisted, though all the important calculations are done with exact integer arithmetic. The reader can download the code for Theorems 1.1 and 1.2 respectively at
[…]
@@ 1.7 Computer Assistance
I used ChatGPT-4o and ChatGPT-5 in various ways to help me with this project. Since this is my first time doing a project with the assistance of ChatGPT, I think that this warrants some discussion. I used ChatGPT to discuss high level mathematical ideas, to find typos and other inconsistencies in my paper, to format data both for the paper and for my computer code, and also to document the efficiency of my programs. I did not let ChatGPT compose prose for me and I ignored its writing advice. I want to retain $100$ percent of my own voice when I write something. Overall, I found ChatGPT to be a helpful and enthusiastic companion as I worked.
@@ 1.8 Acknowledgements
I thank Samuel Lelievre and Alba Malaga for telling me all about flat tori (over a period of some years) and in particular telling me about the flatness problem and supplying me with some historical context. The main thing that inspired me to work on this project was the infectious enthusiasm of Samuel and Alba. I also thank Peter Doyle, Fabian Lander, Noah Montgomery, Stepan Paul, and Saul Schleimer for helpful discussions about this material. I thank Frank Lutz for making his manifolds webpage, which lets you download the combinatorial data for triangulations of surfaces and other manifolds. This page was very useful to me.
@@ 2.2 Proof of the Main Theorem
This completes the proof of the Main Theorem modulo the Hull Lemma.
@@ 2.3 Proof of the Hull Lemma
2.4 The Hull Theorem
For the interested reader we sketch the proof of the result from [BE] that a $7$-vertex embedded polyhedral torus cannot have all $7$ vertices on its convex hull boundary. We call this result the Hull Theorem. This result is not needed for any of our results.
If such an example $\Omega$ exists then up to combinatorial isomorphism its internal edge pattern must be as in Figure 2.2. We think of $\Omega$ as a subset of projective space $\mbox{\boldmath{$P$}}^{3}$. We can apply a projective transformation so that vertex $(2)$ moves to the point $[0:0:1:0]$ at infinity in $\mbox{\boldmath{$P$}}^{3}$. This point is “infinitely far away” along the $Z$-axis. We also can arrange that the (now) rays $(2j)$ start at $(j)$ and move downward (rather than upwards) along the $Z$-axis, limiting on $(2)$. We do all this by a projective transformation that maps $H-\{(2)\}$ into the affine patch $\mbox{\boldmath{$R$}}^{3}$.
Now, $H-\{(2)\}$ is still a convex subset in $\mbox{\boldmath{$R$}}^{3}$. It is like a hexagonal prism that has been truncated at one end. Here is the crucial observation: since $H$ is convex, the projection of the hexagon $(603541)$ into the $XY$-plane is a convex hexagon. We can further normalize so that the projections of $(15)$ and $(36)$ into the $XY$-plane are parallel line segments. We do this by mapping the line $215]\cap[236]$ to a line at infinity which contains $(2)$. Here $[abc]$ is the plane containing $(a)$ and $(b)$ and $(c)$.
Figure 2.3: Projection of part of $\partial H$ into the $XY$-plane
The only triangles of $\partial H$ not in $\Omega$ are $(036)$ and $(145)$. By convexity, these two triangles are bent downward. What we mean is that the plane containing $(036)$ has the rest of $H$ beneath it. The same goes for the plane containing $(145)$.
The blue triangle in Figure 2.3 is the internal triangle $(346)$. We can foliate this triangle by parallel line segments as shown in the figure. These segments are parallel in space and they project to parallel segments in $\mbox{\boldmath{$R$}}^{2}$. The red triangle in Figure 2.3 is the internal triangle $(015)$. We make all the same constructions for this triangle.
Say that a special planes $\mbox{\boldmath{$R$}}^{3}$ whose projections to $\mbox{\boldmath{$R$}}^{2}$ is a line parallel to the projections of our red and blue foliations. One special plane contains $(15)$. In this plane, the red foliation is above the blue foliation. Another special plane contains $(36)$. In this plane, the blue foliation is above the red foliation. So, by the intermediate value theorem, there is a special plane for which these two foliations coincide. But then $(015)$ and $(346)$ intersect. This contradicts the fact that $\Omega$ is embedded.
3 Robust Embedding
3.1 Main Result
Let $T=T(z_{0},z_{1},z_{2})$ be the example from Equation 1. Let $T(z^{\prime}_{0},z^{\prime}_{1},z^{\prime}_{2})$ be as in Equation 1, but with variables $z_{0}^{\prime},z_{1}^{\prime},z_{2}^{\prime}$. We do not insist that $z_{0}^{\prime},z_{1}^{\prime},z_{2}^{\prime}$ are integers. Let
$|T^{\prime}-T|=\max(|z_{0}^{\prime}-z_{0}|,|z_{1}^{\prime}-z_{1}|,|z_{2}^{\prime}-z_{2}|).$ (5)
$T$ is $K$-robustly embedded if and only if $T^{\prime}$ is embedded whenever $|T^{\prime}-T|\leq K$. Here is our main result.

### v9

@@ 1.1 Context and Main Results
A flat torus is a quotient of the form $\mbox{{$R$}}^{2}/\Lambda$, where $\Lambda$ is a lattice of translations of $\mbox{{$R$}}^{2}$. More concretely, a flat torus is what you get when you identify the opposite sides of a parallelogram by translations.
A polyhedral torus is a continuous, piecewise affine embedding $\phi:T\to\Omega\subset\mbox{{$R$}}^{3}.$ Here $T$ is a flat torus that has been triangulated, and $\phi$ is a continuous embedding that is affine on each triangle of the triangulation. If, additionally, $\phi$ is an affine isometry on each triangle, we say that $\phi$ is flat. When $\phi$ is a flat embedding, $\phi$ tells how to build a torus in $\mbox{{$R$}}^{3}$ out of finitely many triangles so that the cone angle around each vertex is $2\pi$. We call a flat embedded polyhedral torus a paper torus. Some authors would call this an origami torus.
Surprisingly, paper tori exist. The 1960 paper of Y. Burago and V. Zalgaller [BZ1] gives the first constructions. Some of the simplest paper tori are called diplotori. These are described by U. Brehm [Br] in 1978, and also described in H. Segerman’s book [Se, §6].
The 1995 paper [BZ2] proves that one can realize every isometry class of flat torus as a paper torus. The works of T. Tsuboi [T], and (independently) P. Arnoux, S. Lelièvre, and A. Málaga [ALM], show that every flat torus, except those made by identifying the sides of a rectangle, is realized by some diplotorus. (Both papers have a separate construction which works for the rectangular flat tori.) The 2024 preprint of F. Lazarus and F. Tallerie [LT] gives a universal combinatorial type of triangulation which does the job simultaneously for all isometry types. Their triangulation $1217$ vertices (and $2434$ faces).
There are diplotori, based on a pair of regular pentagons, which have $10$ vertices. In 2025, Vincent Tugayé [Tu], discovered a $9$-vertex paper torus. At the same time, one needs at least $7$ vertices to make a paper torus because there is no $6$ vertex triangulation of a torus.
The $7$-vertex triangulation of a torus, the Moebius triangulation, is unique up to combinatorial isomorphism. The Moebius triangulation gives the famous embedding of the complete graph $K_{7}$ in a torus. Figure 1.1 shows part of the universal cover of the Moebius torus, as well as a fundamental domain for it.
@@
1.2 Seven Vertices
@@
In order to keep this paper self-contained, I will give a light and entirely combinatorial computer-assisted proof of the Hull Lemma. It is not very hard to promote the Hull Lemma to the Hull Theorem, but since this is not needed for the proof of Theorem 1.1 I will not do it. I give the argument in my supplementary notes [S].
1.3 Eight Vertices
@@
There are $7$ combinatorial types of triangulation of the torus having $8$ vertices. Of these $7$ there is (in my opinion) a worst triangulation and a best triangulation. The worst triangulation is obtained from the Moebius triangulation by subdividing one of the triangles. The best triangulation has degree sequence $66666666$ and is vertex-transitive. This triangulation is very much like the Moebius triangulation. Figure 1.2 below shows part of the universal cover of this triangulation, as well as a fundamental domain. I will explain the meaning of the blue triangles below. The triangles all have the form $(a,a+1,a+3)$ and $(a,a+2,a+3)$ with indices taken mod $8$.
@@
After an intense period of experimentation, I discovered some $8$-vertex paper tori. In §4 I will explain in detail how I found these. In the example below all the coordinates except $z_{0},z_{1},z_{2}$ are exact. The listings of $z_{0},z_{1},z_{2}$ are $32$-digit truncations of the true values.
[…]
Following a suggestion of Peter Doyle, I call this example and ones like it pup tents, on account of their appearance. See Figure 4.1. Specifically, a pup tent is an $8$-vertex paper torus with the following properties:
  • It is based on the $8$-vertex uniform triangulation.
  • It has order $2$ symmetry given by the map $(x,y,z)\to(-x,-y,z)$.
  • It has $6$ faces in its convex hull boundary, corresponding to the blue faces in Figure 1.2.
Here I discuss the proof that the example in Equation 1 is really the truncation of an $8$-vertex paper torus. If we leave $z_{1},z_{2},z_{3}$ unspecified, the example in Equation 1 is part of a $3$-parameter family of examples we call $T(z_{1},z_{2},z_{3})$. Let $\theta_{k}(z_{1},z_{2},z_{3})$ be the cone angle of $T(z_{1},z_{2},z_{3})$ at vertex $k$. Let
$F(z_{1},z_{2},z_{3})=(\theta_{1},\theta_{2},\theta_{3}).$ (2)
If $F(z_{1},z_{2},z_{3})=(2\pi,2\pi,2\pi)$ and $T(z_{1},z_{2},z_{3})$ is embedded then we have our proof. Let $dF$ be the Jacobian of $F$, computed at our example. Here we show $dF$ with $4$ digits of accuracy.
$\left[\begin{matrix}-1.6627\ldots&1.4699\ldots&-0.0244\ldots\\ 1.4699\ldots&0.04714\ldots&-1.1300\ldots\\ -0.0244\ldots&-1.1300\ldots&1.9196\ldots\end{matrix}\right]$ (3)
The matrix $dF$ is an invertible matrix, and all of its eigenvalues exceed $3/10$ in absolute value. This means that the inverse $dF^{-1}$ is also small. In other words, $dF$ is a well-conditioned matrix.
Our proof is an interplay of three things. First, we show that our example is robustly embedded in the sense that moderate perturbations keep it embedded. Second, as we say above, our example is very close to being flat. Third we show, by bounding derivatives, that our Jacobian matrix stays fairly constant and well conditioned throughout a healthy neighborhood of our example. We then use an effective version of the Inverse Function Theorem so show that a true paper torus is quite near our example. We will add enough slack to our proof so that it would work for examples which are fairly near the one in Equation 1.
Given the flexibility of our argument, we get the following easy corollary.
Corollary 1.4
The space $\cal X$ of pup tents contains an open $6$-dimensional ball.
Proof: We start with the example based on Equation 1 and we then perturb our points symmetrically by changing the $X$ and $Y$ coordinates but not the $Z$-coordinates. As long as the perturbation is small enough, the Inverse Function Theorem selects $Z$-coordinates for us which give a pup tent. We have $8$ free parameters to play with and then these are grouped into a $2$-dimensional foliation under the action of similarities of the $XY$-plane which fix the origin. A cross-section of this foliation gives us our $6$-dimensional open manifold of inequivalent pup tents. $\spadesuit$
1.4 Further Developments
Since I wrote the first version of this paper, there have been a number of developments. First of all, some of us at an ICERM workshop in August 2025 printed out pup tents from a paper model I made (with advice from Alba Málaga). With some amount of dexterity, you can use the paper model in [S] to make your own. This model is based on a pup tent that has slightly different coordinates than the one in Equation 1, but the shape is so close you can hardly notice the difference.
Second, Peter Doyle made an amazing discovery of some degenerate versions of the pup tent which realize every flat structure. We call the set of these degenerate pup tents the golden valley. Peter and I then used the golden valley to get some further resuts about $8$ vertex paper tori. Our sequel paper [DS] has many of our findings. Here are two example results:
@@ 1.4 Further Developments
    Any flat torus which lacks reflection symmetry – this accounts for almost all of them – is realized by a pup tent.
@@
    Given any $\epsilon>0$ there is an $8$-vertex paper torus within $\epsilon$ of the unit equilateral triangle in the Hausdorff metric.
Also, the proofs in [DS] give a second and independent proof of Theorem 1.2.
Third, after hearing about the pup tent, my undergraduate student, Zhengyu (Byron) Zou tried for similar results in hyperbolic geometry. Recently he found a $10$-vertex intrinsically hyperbolic embedding of a genus $2$ surface in $\mbox{{$H$}}^{3}$. This is vertex-minimal, because you need $10$ vertices to triangulate a genus $2$ surface. See [Z]. Even more recently, he discovered a $12$-vertex hyperbolic embedding of a genus $2$ surface which has order $2$ symmetry. The order $2$ symmetry implements the hyperelliptic involution.
1.5 Computer Assistance
@@
I used ChatGPT-4o and ChatGPT-5 in various ways to help me with this project. Since this is my first time doing a project with the assistance of ChatGPT, I think that this warrants some discussion. I used ChatGPT to discuss high level mathematical ideas, to find typos and other inconsistencies in my paper, to format data both for the paper and for my computer code, and also to document the efficiency of my programs. There was one astounding thing ChatGPT5 did. In a stroke of brilliance, it found on its own the expression for the denominator of $g_{11}$ at the end of the paper. I might not have arrived at that. I did not let ChatGPT compose prose for me and I ignored its writing advice. I want to retain $100$ percent of my own voice when I write something. Overall, I found ChatGPT to be a helpful and enthusiastic companion as I worked.
@@ 1.6 Acknowledgements
I thank Samuel Lelièvre and Alba Málaga for telling me all about flat tori (over a period of some years) and in particular telling me about the flatness problem and supplying me with some historical context. The main thing that inspired me to work on this project was the infectious enthusiasm of Samuel and Alba. I also thank Peter Doyle, Fabian Lander, Noah Montgomery, Stepan Paul, Saul Schleimer, and Byron Zou for helpful discussions about this material. I thank Frank Lutz for making his manifolds webpage, which lets you download the combinatorial data for triangulations of surfaces and other manifolds. This page was very useful to me.
@@ 2.2 Proof of the Main Theorem
This completes the proof of the Main Theorem modulo the Hull Lemma. We mention again that the Hull Lemma is an immediate consequence of (what I call) the Hull Theorem in [BE]. Readers who are keen to see the existence of the $8$-vertex paper torus can stop reading now. Readers who want a self-contained proof of the Hull Lemma should keep reading.
@@ 2.3 Proof of the Hull Lemma
The following lemma lets us deduce the cycle structure on $H$ from the pattern of internal edges.
@@
Remark: As mentioned in the introduction, we have a separate geometric argument that rules out the $6$ patterns like the one in Figure 2.2, thereby giving a second proof of the Hull Theorem. See my supplementary notes [S] for this argument.
3 Proof of Theorem 1.2
Throughout this chapter $\|\cdot\|$ denotes the Euclidean norm and $\|\cdot\|_{\infty}$ denotes the maximum of the absolute values of vector or matrix entries. Also, $\angle$ denotes angles.
3.1 Robust Embedding
Let $T=T(z_{1},z_{2},z_{3})$ be the example from Equation 1. Let $T(z^{\prime}_{1},z^{\prime}_{2},z^{\prime}_{3})$ be as in Equation 1, but with variables $z_{1}^{\prime},z_{2}^{\prime},z_{3}^{\prime}$. Let
$|T^{\prime}-T|=\max(|z_{1}^{\prime}-z_{1}|,|z_{2}^{\prime}-z_{2}|,|z_{3}^{\prime}-z_{3}|).$ (5)
We say that $T$ is $K$-robustly embedded if and only if $T^{\prime}$ is embedded whenever $|T^{\prime}-T|\leq K$. In this section we prove the following result.